OpenStudy (anonymous):

1/3 2c-5 +3=7 i really need some hlep on this one @ZeHanz

OpenStudy (zehanz):

YOur equation is not entirely clear to me...

OpenStudy (anonymous):

whats 1/3 divided by 4

OpenStudy (zehanz):

1/12

OpenStudy (anonymous):

whis 5+1/12

OpenStudy (zehanz):

Because\[\frac{ \frac{ a }{ b } }{ c }=\frac{ a }{ b }\cdot \frac{ 1 }{ c }=\frac{ a }{ bc }\]

OpenStudy (zehanz):

5 is 60/12 so now you can add them up...

OpenStudy (anonymous):

whis 14 dvied by 7/4